OpenStudy (anonymous):

A basketball drops from a height of 76 in. On bounce 1, it rebounds to a height of 38 in. On bounce 2, it rebounds to a height of 19 in. How high does the basketball bounce on bounce 3? A.9in B.19in C.9.5in D.4.5in

OpenStudy (anonymous):

Pattern: 76/2 = 38in, 38/2 = 19in ⇒ 19/2 = 9.5in = answer C